OpenStudy (anonymous):

factorise by splitting into two's: df+6g-6f-dy i got this: = (6y-6f) + (df-dg) = 6(g-f) +d(f-g) this is wrong, why?

OpenStudy (cwrw238):

you have a dy in the question and dg in your answer - have you got question right?

OpenStudy (cwrw238):

if the question contains dg df+6g-6f-dg = 6g - 6f + df - dg you need to have g-f in both parentheses = 6(g-f) - d(g-f) = (6-d)(g-f)

OpenStudy (cwrw238):

ok ? - i think it must be dg in original question

OpenStudy (anonymous):

yeah it was a dg the answer you gave us wrong though. the answer in book is (f-g) (d-6)

OpenStudy (cwrw238):

ok lets expand (f-g) (d-6) = df - 6f - dg + 6g - which is the same as the question except terms are moved around a bit - so the book is correct now try my answer (6-d)(g-f) = 6g - 6f -dg + df which is also same as the question both answers are correct - its just that the order of terms in each bracket has changed

OpenStudy (cwrw238):

do you follow me?

OpenStudy (anonymous):

oh i get it now thx.

OpenStudy (cwrw238):

yw